OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3).

OpenStudy (anonymous):

2x-y=7

OpenStudy (anonymous):

Parallel lines must have the same gradient. The gradient of a line is m in the equation y=mx+c. So the gradient must be 2. Now by using the general formula for lines y-y1=M(x-x1) you can put in the coordinates (5.3) to get , y-3=2(x-5) which is the same as y=2x-7.

OpenStudy (anonymous):

So it's y = 2x - 7.

OpenStudy (anonymous):

Hey, thank you! Would you mind answering a few more?

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2).

OpenStudy (anonymous):

Perpendicular lines have the negative reciprical of the original line. So you basically take the orginal gradient and put it under 1. So the perpendicular gradient to 2 would be -1/2. Then just put that into the general formula. y-y1=M(x-x1) where y1 is 3 and x1 is 6.

OpenStudy (anonymous):

For the next one, get the formula to look like y=mx+c. So y=3x+3. Then use the gradient of 3 and put it into the formula.

OpenStudy (anonymous):

And, I also have Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2). Explain, in complete sentences, the relationship between the equations of perpendicular lines.